OpenStudy (anonymous):

The speed of sound V near the Earth's surface can be represented by this equation V=20 radical t+273, where t is the temperature in degrees Celsius. Find the temperature when V=380 meters per second. Round to the nearest whole number. give me one sec and i will show you my problem

OpenStudy (anonymous):

\[V=20\sqrt{t+273}\]

OpenStudy (anonymous):

\[380=20\sqrt{(t+273)}\] \[19=\sqrt{(t+273)}\] \[19^{2}=(t+273)\] 361-273= t t = 88 C

OpenStudy (anonymous):

got the same

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

ur welcome. and #krypton, thanks for the affirmation.

OpenStudy (anonymous):

anytime :)